Can someone please help me with the following statistic question?
(Hypothesis Test Summary) (Show all work and steps)
4) According to a U.S. Census Bureau survey during 2005-2007, the daily one-way commute time of U.S. workers averages 25 minutes with , we'll assume, a standard deviation of 13 minutes. An investigator wishes to determine whether the national average describes the commuter time for all workers in the Chicago area. Commute times are obtained for a random sample of 169 workers from this area, and the mean time is found to be 22.5 minutes. Test the Null Hypothesis at the .05 level of significance

Answer by stanbon(75887)

According to a U.S. Census Bureau survey during 2005-2007, the daily one-way commute time of U.S. workers averages 25 minutes with , we'll assume, a standard deviation of 13 minutes.
An investigator wishes to determine whether the national average describes the commuter time for all workers in the Chicago area.
Commute times are obtained for a random sample of 169 workers from this area, and the mean time is found to be 22.5 minutes.
Test the Null Hypothesis at the .05 level of significance
----
Ho: u = 25
Ha: u # 25
----
x-bar = 22.5
z(22.5) = (22.5-25)/[13/sqrt(169)] = -2.5
-----
p-value = 2*P(z < -2.5) = 2(normalcdf(-100,-2.5) = 0.0124
-----
Since the p-value is less than 5%, reject Ho
Conclusion: 25 is not the average for all workers in te Chicago area.
================
Cheers,
Stan H.